OpenStudy (anonymous):

given: x=2 y=0 find (r,theta)

OpenStudy (anonymous):

r= \[\sqrt{x ^{2}+y ^{2}} = \sqrt{4}=2\] \[\theta =?\]

hartnn (hartnn):

x=r cos theta y=r sin theta theta =tan inverse of (y/x)

OpenStudy (anonymous):

but what if i don't know the value of tan inverse @hartnn because we arent allowed calculators in the exam

hartnn (hartnn):

here tan inverse of 0 is 0----> this is easy but you have r value so u can use x=r cos theta or y=r sin theta to get theta,if u can find sin inverse or cos inverse without calcy atleast 1 of sin or cos or tan inverse is must for theta

OpenStudy (anonymous):

ah okay, thank you for telling me about tan inverse zero, makes working easier

OpenStudy (anonymous):

so \[\Theta = 0\]

hartnn (hartnn):

yes,theta = 0